# You Only Forward Once: An Efficient Compositional Judging Paradigm

### v2

@@ 1 Introduction
Large language models (LLMs) [team2024qwen2, yang2025qwen3, ahmed2025qwen, bai2023qwen, liu2024deepseek, guo2025deepseek] have achieved remarkable progress, and advances in multimodal learning now enable them to process diverse modalities. Multimodal Large Language Models (MLLMs) [bai2025qwen2, wang2024qwen2, bai2023qwenvl, lu2025ovis2, lu2024ovis] have recently demonstrated significant promise in information matching due to their ability to jointly process images, videos, and natural language queries, making them far more applicable in real-world scenarios than traditional unimodal retrievers. In particular, state-of-the-art MLLM-based rerankers harness powerful vision encoders and causal attention mechanisms to effectively fuse visual and textual information, usually ultimately regressing a single relevance score, and achieves strong performance in cross-modal matching tasks such as image search and e-commerce recommendation. However, despite these advances, outputting only a scalar relevance score leads to substantial information loss, as fine-grained semantic distinctions in complex user queries—such as compositional intent, attribute-level preferences, or contextual references—are inevitably collapsed into a single value. As an example, Fig. 1 illustrates a recommendation failure of the state-of-the-art (SOTA) multimodal reranker, Jina-Reranker-M0 [jina2025blog], which prioritizes superficial semantic overlap between the user’s query and the pink dress, resulting in an inappropriately high ranking for this item. Therefore, achieving accurate and fine-grained understanding of intricate multimodal queries has become an urgent challenge that must be addressed to realize the full potential of next-generation information matching systems.
@@
(a) Representation-based Model
(b) Adapted MLLM for Scoring
(c) Autoregressive Model
(d) Ours: YOFO
@@
Recent studies on matching can be grouped into four categories. 1) Representation-based architectures, as illustrated in Fig. 2(a), independently embed the query and the document and then compute their similarity [radford2021learning, huang2013learning, gao2014modeling, shen2014learning, palangi2016deep, wan2016deep, cohen2016end]. However, fixed-size embeddings limit these models’ capacity, making it difficult to capture all relevant information and to meet fine-grained requirements [xu2025survey]. 2) Interaction-based models have emerged to mitigate these limitations: they first compute low-level interaction representations between the query and document, and then a neural network analyzes these patterns to output a scalar relevance score for the query–document pair [hu2014convolutional, dai2019deeper, nogueira2019multi, gao2022long]. These approaches achieve higher ranking quality than representation-based methods but are computationally more expensive and remain constrained by representational dimensionality. 3) MLLM/LLM-based architectures, adapted to output a single relevance score improve matching quality by leveraging the deep semantic understanding and long-context capacity of large language models [ma2024fine]. However, this design is misaligned with the native generative objective of LLMs and undermines fine-grained input understanding. 4) LLMs and MLLMs can serve as native judges in a paradigm termed generative matching [yoon2024listt5, yang2025rank], which reframes the task as autoregressive generation aligned with their generative nature. Nevertheless, these methods incur high computational overhead due to token-by-token decoding, making them difficult to deploy in real-world applications.
@@
To evaluate the efficacy of our method, we derive a training set from SA-1B [kirillov2023segment] and train the model on it to improve generalizability. We then construct a test set from LRVS-Fashion [lepage2023lrvs] to assess recommendation performance in a different domain. Extensive experiments demonstrate that our method accurately produces judgments conditioned on the input and template, and that later judgments can condition on earlier ones, enabling dependency-aware analysis. Furthermore, YOFO also benefits from post-hoc Chain-of-Thought (CoT), which encourages the model to reason implicitly before judging.
@@ 2.1 Multimodal Large Language Models
Multimodal Large Language Models (MLLMs) [bai2025qwen2, wang2024qwen2, bai2023qwenvl, lu2025ovis2, lu2024ovis, hong2025worldsense, ma2024visual, ma2024multi, xue2023clip] have significantly extended the capabilities of Large Language Models (LLMs) [team2024qwen2, yang2025qwen3, ahmed2025qwen, bai2023qwen, liu2024deepseek, guo2025deepseek, li2025adaptive] by enabling them to perceive, reason over, and generate responses grounded in multimodal inputs. However, the dense visual token sequences required for such alignment incur prohibitive computational costs. Recent approaches address this by either compressing tokens while preserving fidelity or rethinking how visual information is integrated. For instance, CrossLMM [yan2025crosslmm] employs dual cross-attention mechanisms to drastically reduce token count with minimal performance degradation. Similarly, MM-GEM [ma2024multi] uses a PoolAggregator to support both generative and embedding tasks. These advances demonstrate that strategic token reduction can maintain multimodal performance at lower cost.
@@ 2.2 Rerankers
Reranking serves as a critical refinement stage in modern retrieval systems, aiming to re-order an initial candidate set by modeling fine-grained query–document interactions. Early neural rerankers employed representation-based architectures, calculating embeddings independently and then similarity, such as DSSM [huang2013learning, gao2014modeling], Convolutional DSSM [shen2014learning]) and Long Short-Term Memory (LSTM) variants [hochreiter1997long, palangi2016deep, wan2016deep, cohen2016end]. But this kind of models are limited by the fixed-size embedding, and thus were soon superseded by interaction-based cross-encoders, most notably BERT-based models, that concatenate query and document as input and predict relevance via a classification or regression head over the [CLS] token. [nogueira2019multi] first introduced this method and proved it effective. These models achieve strong performance but require pairwise inference, limiting scalability, and is constrained by fixed representational dimention and content length.   With the rise of LLMs, researchers have repurposed powerful encoder-decoder (e.g., T5 [raffel2020exploring]) or decoder-only (e.g., LLaMA [touvron2023llama]) backbones for reranking. A dominant paradigm [zhuang2023rankt5, ma2024fine] fine-tunes LLMs to regress continuous relevance scores or classify relevance levels, leveraging their rich contextual understanding and long-context capabilities for improved semantic matching. Jina-Reranker-M0 also follows this paradigm but leverages Qwen2-VL [wang2024qwen2], enabling support for multimodal inputs.   More recently, generative reranking has emerged as a paradigm shift: instead of scoring candidates individually, the model directly generates the ranked sequence—such as a list of document indices or IDs—conditioned on the full candidate set. Methods like [yoon2024listt5, yang2025rank] process all candidates jointly and autoregressively output the optimal ranking. This generative formulation naturally integrates reranking into end-to-end AI systems and shows particular promise in multimodal and agent-driven retrieval scenarios.
@@ 3 Method
In this section, we introduce YOFO, an efficient judging paradigm adapting an autoregressive model to receive a series of requirements as input and yields the judgements (“yes” or “no”) for all requirements, respectively, as output. We begin by formulating the problem in Sec. 3.1, followed by the overall architecture of YOFO in Sec. 3.2, training in Sec. 3.3, inference in Sec. 3.4.
(a) YOFO Training
(b) YOFO Inference
@@ 3.1 Problem formulation
As illustrated in Sec. 1, models akin to Jina-Reranker-M0 [jina2025blog] struggle to perform accurate cross-modal retrieval. One might attribute this to visual information degradation caused by the image encoder and the connector projection, which impairs fine-grained alignment. However, we empirically find that these models also fail to retrieve text accurately from natural-language queries. As shown in Tab. 1(a), Jina-Reranker-M0 assigns a higher score to winter trousers simply because they are gray, failing to recognize that the query explicitly requires summer-appropriate clothing, while “gray” is only a conditional preference. After further simplifying the prompt, Tab. 1(b) shows that Jina-Reranker-M0 fails to handle negation reliably, appearing to measure lexical overlap rather than semantic similarity on slightly more complex inputs. Verifying each requirement in the query individually would likely yield better results. To this end, we formally define the compositional judging problem.
query documents document 1 document 2
query 0.88 0.81
(a) Qeery: “I’m looking for some bottoms that are ideally suitable for summer wear. And I’d prefer them to be grey, if available.”, document 1: “Grey formal trousers, a touch of warmth—perfect for winter.”, document 2: “Black skirt, perfect for casual or date outfits in summer.”
query documents document 1 document 2
query 0.82 0.94
(b) Qeery: “I’m looking for a long-sleeve without a logo on the chest.”, document 1: “A long-sleeve hooded.”, document 2: “A long-sleeve with a fairly large logo on the chest.”
Table 1: Relevance scores computed by jina-reranker-m0 on text-to-text retrieval examples.
@@ 3.2 Overall Architecture
YOFO leverages a decoder-only MLLM as the backbone. As illustrated in Fig. 3, we add a single token unknown to the end of each requirement $\bm{p}_{i}$, and concatenate all requirements to form the template $\bm{t}$. The template $\bm{t}$ is then tokenized into $\bm{s}$, and the indices of the single token unknown in the sequence are denoted $(pos_{i})_{i=1}^{N}$. After a single forward pass, the model must predict whether each unknown position should be “yes” or “no”. The forward process can be formulated as:
@@ 3.3 Training
Without Post-hoc CoT. As shown in Fig. 3(a), we apply supervision at the answer positions to train the LLM to judge the requirements correctly. Specifically, we minimize the cross-entropy loss between the model logits $(\bm{l}_{i})_{i=1}^{N}$ and the ground-truth labels. The answer loss $\mathcal{L}_{\text{answer}}$ is defined as:
@@
Although this objective resembles the standard autoregressive loss, when the model computes $\bm{l}_{i}$ for the $i$-th requirement, it does not observe the answers to previous requirements. Therefore, $\mathcal{L}_{\text{answer}}$ fundamentally differs from the autoregressive objective. Optimizing $\mathcal{L}_{\text{answer}}$ enables the LLM to make dependency-aware judgments, allowing later decisions to condition on its earlier judgements, as shown in Sec. 4.5.
With Post-hoc CoT. As shown in Fig. 3(a), YOFO with Post-hoc CoT applies supervision to both the answers and the accompanying rationales. The requirements themselves are not supervised, as they are provided as input rather than predicted. Let $\mathcal{V}$ denote the positions of tokens in the reason field, that is $\mathcal{V}=\{i|s_{i}\text{ lies in the reason field}\}$. For each $i\in\mathcal{V}$, the logits at the preceding position (i1)(i-1) predict the distribution over the next token $s_{i}$ in the reason field. Accordingly, we define the following autoregressive loss to encourage the model to consider rationales prior to making judgments:
@@ 3.4 Inference
Fig. 3(b) illustrates how a YOFO-trained LLM performs inference. At inference time, there are no ground-truth annotations for rationales; consequently, the rationale field is omitted from the template. The MLLM backbone ingests the image $I$ and the template $\bm{t}$ and performs a single forward pass, after which the logits $(\bm{l}_{i})_{i=1}^{N}$ can be extracted as in Eq. 3. Finally, we apply the binary decision rule in Eq. 5 to ensure that the LLM outputs only “yes” or “no”. The resulting judgments can then be consumed by downstream tasks—for example, provided as input to an agent for further analysis or mapped to a single score under a predefined criterion.
As an illustrative use case of YOFO, Fig. 3(b) depicts how it can be integrated into rerankers. To obtain a structured template, an LLM first decomposes the user’s query into a set of requirements. These requirements are then assembled into a template and passed, together with the image, to the MLLM. Finally, all requirements are judged in a single forward pass, yielding results that substantially improve reranking quality. Task-specific rules can be devised to exploit these judgments; we leave the design open to practitioners.
@@ 4 Experiments
In this section, we present extensive experiments and ablation studies to validate our approach. We first compare YOFO with Jina-Reranker-M0 [jina2025blog], a representative state-of-the-art (SOTA) multimodal reranker, on the test set derived from LAION-RVS-Fashion [lepage2023lrvs]. We then assess YOFO’s dependency-aware judgement, the effectiveness of post-hoc CoT, and the contribution of individual architectural components.
4.1 Impementation Details
We build on Qwen2-VL-2B-Instruct [wang2024qwen2] and Qwen3-VL-2B-Instruct [qwen3vl2025blog], freezing the vision encoder during training. We fine-tune the LLM using LoRA [hu2022lora]. To conveniently locate the unknown token, we add four special tokens to the Qwen2-VL-2B-Instruct and Qwen3-VL-2B-Instruct processor: <|auth_start|>, <|auth_end|>, <|reason_start|>, and <|reason_end|>. The first two mark the start and end of the answer span, and the last two mark the boundaries of the reason span. Consequently, these special tokens enable straightforward label masking, allowing supervision to be applied exclusively to answers and reasons.
Training is conducted for one epoch with a learning rate of $1\times 10^{-4}$, using the AdamW optimizer. We employ a cosine learning rate scheduler with a warmup ratio of 0.05. We set the maximum sequence length to 1,536 tokens. We use DeepSpeed ZeRO-3 for distributed training. We load the training data with eight worker processes, dropping the last incomplete batch. Training runs on eight H100 GPUs over four hours. We use bfloat16 mixed precision and enable gradient checkpointing to reduce memory usage. Moreover, we adopt $\lambda=0.55$ as the default setting in all experiments using post-hoc CoT.
(a) Training set example
(b) Test set example
Figure 4: Examples of the training and test sets. Each training sample consists of an image, a set of properties, answers, and corresponding reasons. Each test sample consists of two images, a customer-style query, the corresponding Python variables, ground-truth labels, and the expression used to compute each image’s final recommendation score.
(a) Prompt to construct the training set
(b) Prompt to construct the test set
Figure 5: Prompts for dataset construction. (a) We prompt an MLLM to propose ten properties of a randomly selected image, with an equal number of satisfied and unsatisfied properties. (b) We prompt an MLLM to generate a customer-style query for an image pair, indicate whether the first image satisfies the query better than the second, decompose the query into a set of requirements, and produce an expression used to compute each image’s final recommendation score during YOFO evaluation.

### v3

@@ Abstract
Multimodal Large Language Models, Cross-modal Retrieval, MLLM-as-Judge
@@ 1 Introduction
Large language models (LLMs) (Team and others, 2024; Yang et al., 2025a; Ahmed et al., 2025; Bai et al., 2023a; Liu et al., 2024; Guo et al., 2025) have achieved remarkable progress, and advances in multimodal learning now enable them to process diverse modalities. Multimodal Large Language Models (MLLMs) (Bai et al., 2025b; Wang et al., 2024; Bai et al., 2023b; Lu et al., 2025, 2024) have recently demonstrated significant promise in information matching due to their ability to jointly process images, videos, and natural language queries, making them far more applicable in real-world scenarios than traditional unimodal retrievers. In particular, state-of-the-art MLLM-based rerankers harness powerful vision encoders and causal attention mechanisms to effectively fuse visual and textual information, usually ultimately regressing a single relevance score, and achieves strong performance in cross-modal matching tasks such as image search and e-commerce recommendation. However, despite these advances, outputting only a scalar relevance score leads to substantial information loss, as fine-grained semantic distinctions in complex user queries—such as compositional intent, attribute-level preferences, or contextual references—are inevitably collapsed into a single value. As an example, Figure 1 illustrates a recommendation failure of the state-of-the-art (SOTA) multimodal reranker, Jina-Reranker-M0 (Jina AI, 2025), which prioritizes superficial semantic overlap between the user’s query and the pink dress, resulting in an inappropriately high ranking for this item. Therefore, achieving accurate and fine-grained understanding of intricate multimodal queries has become an urgent challenge that must be addressed to realize the full potential of next-generation information matching systems.
Recent studies on matching can be grouped into four categories. 1) Representation-based architectures, as illustrated in Figure 2 (a), independently embed the query and the document and then compute their similarity (Radford et al., 2021; Huang et al., 2013; Gao et al., 2014; Shen et al., 2014; Palangi et al., 2016; Wan et al., 2016; Cohen and Croft, 2016). However, fixed-size embeddings limit these models’ capacity, making it difficult to capture all relevant information and to meet fine-grained requirements (Xu et al., 2025). 2) Interaction-based models have emerged to mitigate these limitations: they first compute low-level interaction representations between the query and document, and then a neural network analyzes these patterns to output a scalar relevance score for the query–document pair (Hu et al., 2014; Dai and Callan, 2019; Nogueira et al., 2019; Gao and Callan, 2022). These approaches achieve higher ranking quality than representation-based methods but are computationally more expensive and remain constrained by representational dimensionality. 3) MLLM/LLM-based architectures, adapted to output a single relevance score improve matching quality by leveraging the deep semantic understanding and long-context capacity of large language models (Ma et al., 2024c), as depicted in Figure 2 (b). However, this design is misaligned with the native generative objective of LLMs and undermines fine-grained input understanding. 4) LLMs and MLLMs can serve as native judges in a paradigm termed generative matching (Yoon et al., 2024; Yang et al., 2025b), which reframes the task as autoregressive generation aligned with their generative nature, as demonstrated in Figure 2 (c). Nevertheless, these methods incur high computational overhead due to token-by-token decoding, making them difficult to deploy in real-world applications.
@@
To evaluate the efficacy of our method, we derive a training set from SA-1B (Kirillov et al., 2023) and train the model on it to improve generalizability. We then construct a test set from LRVS-Fashion (Lepage et al., 2023) to assess recommendation performance in a different domain. Extensive experiments demonstrate that our method accurately produces judgments conditioned on the input and template, and that later judgments can condition on earlier ones, enabling dependency-aware analysis. Furthermore, YOFO also benefits from post-hoc Chain-of-Thought (CoT), which encourages the model to reason implicitly before judging.
@@ 2.1 Multimodal Large Language Models
Multimodal Large Language Models (MLLMs) (Bai et al., 2025b; Wang et al., 2024; Bai et al., 2023b; Lu et al., 2025, 2024; Hong et al., 2025; Ma et al., 2024b, a; Xue et al., 2023) have significantly extended the capabilities of Large Language Models (LLMs) (Team and others, 2024; Yang et al., 2025a; Ahmed et al., 2025; Bai et al., 2023a; Liu et al., 2024; Guo et al., 2025; Li et al., 2025) by enabling them to perceive, reason over, and generate responses grounded in multimodal inputs. However, the dense visual token sequences required for such alignment incur prohibitive computational costs. Recent approaches address this by either compressing tokens while preserving fidelity or rethinking how visual information is integrated. For instance, CrossLMM (Yan et al., 2025) employs dual cross-attention mechanisms to drastically reduce token count with minimal performance degradation. Similarly, MM-GEM (Ma et al., 2024a) uses a PoolAggregator to support both generative and embedding tasks. These advances demonstrate that strategic token reduction can maintain multimodal performance at lower cost.
@@ 2.2 Rerankers
Reranking serves as a critical refinement stage in modern retrieval systems, aiming to re-order an initial candidate set by modeling fine-grained query–document interactions. Early neural rerankers employed representation-based architectures, calculating embeddings independently and then similarity, such as DSSM (Huang et al., 2013; Gao et al., 2014), Convolutional DSSM (Shen et al., 2014)) and Long Short-Term Memory (LSTM) variants (Hochreiter and Schmidhuber, 1997; Palangi et al., 2016; Wan et al., 2016; Cohen and Croft, 2016). But this kind of models are limited by the fixed-size embedding, and thus were soon superseded by interaction-based cross-encoders, most notably BERT-based models, that concatenate query and document as input and predict relevance via a classification or regression head over the [CLS] token. (Nogueira et al., 2019) first introduced this method and proved it effective. These models achieve strong performance but require pairwise inference, limiting scalability, and is constrained by fixed representational dimention and content length.   With the rise of LLMs, researchers have repurposed powerful encoder-decoder (e.g., T5 (Raffel et al., 2020)) or decoder-only (e.g., LLaMA (Touvron et al., 2023)) backbones for reranking. A dominant paradigm (Zhuang et al., 2023; Ma et al., 2024c) fine-tunes LLMs to regress continuous relevance scores or classify relevance levels, leveraging their rich contextual understanding and long-context capabilities for improved semantic matching. Jina-Reranker-M0 also follows this paradigm but leverages Qwen2-VL (Wang et al., 2024), enabling support for multimodal inputs.   More recently, generative reranking has emerged as a paradigm shift: instead of scoring candidates individually, the model directly generates the ranked sequence—such as a list of document indices or IDs—conditioned on the full candidate set. Methods like (Yoon et al., 2024; Yang et al., 2025b) process all candidates jointly and autoregressively output the optimal ranking. This generative formulation naturally integrates reranking into end-to-end AI systems and shows particular promise in multimodal and agent-driven retrieval scenarios.
@@ 3 Method
In this section, we introduce YOFO, an efficient judging paradigm adapting an autoregressive model to receive a series of requirements as input and yields the judgements (“yes” or “no”) for all requirements, respectively, as output. We begin by formulating the problem in Section 3.1, followed by the overall architecture of YOFO in Section 3.2, training in Section 3.3, inference in Section 3.4.
@@ 3.1 Problem formulation
As illustrated in Section 1, models akin to Jina-Reranker-M0 (Jina AI, 2025) struggle to perform accurate cross-modal retrieval. One might attribute this to visual information degradation caused by the image encoder and the connector projection, which impairs fine-grained alignment. However, we empirically find that these models also fail to retrieve text accurately from natural-language queries. For example, consider the query “I’m looking for some bottoms that are ideally suitable for summer wear and I’d prefer them to be grey if available”, paired with two documents: (1) “grey formal trousers, a touch of warmth—perfect for winter” and (2) “black skirt, perfect for casual or date outfits in summer”, Jina-Reranker-M0 assigns a higher score to winter trousers simply because they are gray, failing to recognize that the query explicitly requires summer-appropriate clothing, while “gray” is only a conditional preference. Verifying each requirement in the query individually would likely yield better results. To this end, we formally define the compositional judging problem.
@@ 3.2 Overall Architecture
YOFO leverages a decoder-only MLLM as the backbone. As illustrated in Figure 3, we add a single token unknown to the end of each requirement $\bm{p}_{i}$, and concatenate all requirements to form the template $\bm{t}$. The template $\bm{t}$ is then tokenized into $\bm{s}$, and the indices of the single token unknown in the sequence are denoted $(pos_{i})_{i=1}^{N}$. After a single forward pass, the model must predict whether each unknown position should be “yes” or “no”. The forward process can be formulated as:
@@ 3.3 Training
Without Post-hoc CoT. As shown in Figure 3 (a), we apply supervision at the answer positions to train the LLM to judge the requirements correctly. Specifically, we minimize the cross-entropy loss between the model logits $(\bm{l}_{i})_{i=1}^{N}$ and the ground-truth labels. The answer loss $\mathcal{L}_{\text{answer}}$ is defined as:
@@
Although this objective resembles the standard autoregressive loss, when the model computes $\bm{l}_{i}$ for the $i$-th requirement, it does not observe the answers to previous requirements. Therefore, $\mathcal{L}_{\text{answer}}$ fundamentally differs from the autoregressive objective. Optimizing $\mathcal{L}_{\text{answer}}$ enables the LLM to make dependency-aware judgments, allowing later decisions to condition on its earlier judgements, as shown in Section 4.5.
With Post-hoc CoT. As shown in Figure 3 (a), YOFO with Post-hoc CoT applies supervision to both the answers and the accompanying rationales. The requirements themselves are not supervised, as they are provided as input rather than predicted. Let $\mathcal{V}$ denote the positions of tokens in the reason field, that is $\mathcal{V}=\{i|s_{i}\text{ lies in the reason field}\}$. For each $i\in\mathcal{V}$, the logits at the preceding position (i1)(i-1) predict the distribution over the next token $s_{i}$ in the reason field. Accordingly, we define the following autoregressive loss to encourage the model to consider rationales prior to making judgments:
@@ 3.4 Inference
Figure 3 (b) illustrates how a YOFO-trained LLM performs inference. At inference time, there are no ground-truth annotations for rationales; consequently, the rationale field is omitted from the template. The MLLM backbone ingests the image $I$ and the template $\bm{t}$ and performs a single forward pass, after which the logits $(\bm{l}_{i})_{i=1}^{N}$ can be extracted as in Equation 3. Finally, we apply the binary decision rule in Equation 5 to ensure that the LLM outputs only “yes” or “no”. The resulting judgments can then be consumed by downstream tasks—for example, provided as input to an agent for further analysis or mapped to a single score under a predefined criterion.
As an illustrative use case of YOFO, Figure 3 (b) depicts how it can be integrated into rerankers. To obtain a structured template, an LLM first decomposes the user’s query into a set of requirements. These requirements are then assembled into a template and passed, together with the image, to the MLLM. Finally, all requirements are judged in a single forward pass, yielding results that substantially improve reranking quality. Task-specific rules can be devised to exploit these judgments; we leave the design open to practitioners.
@@ 4 Experiments
In this section, we present extensive experiments and ablation studies to validate our approach. We first compare YOFO with state-of-the-art (SOTA) methods to demonstrate its effectiveness. We then assess YOFO’s dependency-aware judgement, the effectiveness of post-hoc CoT, and the contributions of individual architectural components.
4.1 Implementation Details
We build on Qwen2-VL-2B-Instruct (Wang et al., 2024) and Qwen3-VL-2B-Instruct (Bai et al., 2025a), freezing the vision encoder during training. We fine-tune the LLM using LoRA (Hu et al., 2022). To conveniently locate the unknown token, we add four special tokens to the Qwen2-VL-2B-Instruct and Qwen3-VL-2B-Instruct processor: <|auth_start|>, <|auth_end|>, <|reason_start|>, and <|reason_end|>. The first two mark the start and end of the answer span, and the last two mark the boundaries of the reason span. Consequently, these special tokens enable straightforward label masking, allowing supervision to be applied exclusively to answers and reasons.
Training is conducted for one epoch with a learning rate of $1\times 10^{-4}$, using the AdamW optimizer. We employ a cosine learning rate scheduler with a warmup ratio of 0.05. We set the maximum sequence length to 1,536 tokens. We use DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) for distributed training. We load the training data with eight worker processes, dropping the last incomplete batch. Training runs on eight H100 GPUs over four hours. We use bfloat16 mixed precision and enable gradient checkpointing to reduce memory usage. Moreover, we adopt $\lambda=0.55$ as the default setting in all experiments using post-hoc CoT.
Figure 4: Examples of the training and test sets. Each training sample consists of an image, a set of properties, answers, and corresponding reasons. Each test sample consists of two images, a customer-style query, requirements derived by decomposing the query, ground-truth labels indicating which image better matches the query, and the expression used to compute each image’s final recommendation score.